# LLM-based Agents Suffer from Hallucinations: A Survey of Taxonomy, Methods, and Directions

### v1

@@ 5. Future Directions
Most existing studies investigate hallucination instances and their underlying causes within a single agent loop. However, as we emphasize, agent decision-making is inherently a multi-step and sequential process, in which hallucinations can accumulate and amplify over time. In such cases, hallucinations may initially appear as minor issues, but their iterative accumulation can ultimately lead to severe consequences. Compared to single-step scenarios, hallucination accumulation presents a significantly more complex challenge. Addressing it requires a comprehensive analysis of the agent’s entire decision-making process, thereby facilitating early hallucination detection and mitigation.
Accurate Hallucinatory Location.
@@
As previously mentioned, in contrast to hallucinations in traditional language models that are typically manifested in textual generation errors, agent hallucinations are far more complex, involving full-chain error propagation across multiple interdependent components. While hallucination taxonomy and attribution we presented in Section 3 have mapped these issues to specific modules within the agent, promptly and accurately locate the source of agent hallucinations in the final outputs remains a significant challenge. This difficulty stems from the fact that agent hallucinations may arise at any stage of the decision-making pipeline and often exhibit complex characteristics such as hallucinatory accumulation and inter-module dependency. In fact, the limited amount of research on agent hallucination detection in Section 4.2 and the difficulty of conducting such works are also related to this issue. Therefore, future study should focus on designing agent systems capable of modeling and tracing their entire execution trajectories (Zhang et al., 2025g). For example, lightweight checkpoints can be injected at each stage to verify whether hallucinations have occurred.
Hallucination Mechanistic Interpretability.
@@
Mechanistic interpretability (MI) seeks to uncover how hidden representations and internal components of neural networks give rise to specific behaviors (Olah, 2022; Wang et al., 2022a). Therefore, MI provides a natural pathway for diagnosing and mitigating hallucinations at the root cause. Recent studies demonstrate that MI techniques, such as feature analysis (Ji et al., 2025; Ferrando et al., 2024) and causal pathway tracing (Jin et al., 2024b) can reveal the internal sources of hallucinations in LLMs and provide insights for mitigation. However, extending MI to LLM-based agents introduces new challenges. Unlike LLMs, where interpretability analysis is typically performed under single-step prediction with controlled prompt–response settings, agent hallucinations emerge through multi-step sequential interactions involving reasoning, tool use, memorization, and MAS communication. This dynamic process substantially complicates the controlled interventions required for mechanistic analysis. Future research should therefore adapt MI techniques to account for these dynamic and interconnected processes, enabling more precise diagnosis and systematic mitigation of agent hallucinations.
Unified Benchmark Construction.
@@
Existing benchmarks for agent hallucinations are often limited to a specific hallucination type. For example, Guan et al. (2024) propose HALLUSIONBENCH to evaluate hallucination issues in visual-language reasoning; Hu et al. (2025b) propose MemoryAgentBench to evaluate hallucination issues in agents during memory retrieval and update; Zhang et al. (2024b) introduce ToolBH to evaluate execution hallucinations in tool use across different scenarios from both depth and breadth perspectives. Therefore, there is a lack of a unified benchmark of hallucination evaluation that can define diverse hallucination scenarios and adopt various evaluation metrics to comprehensively assess the extent of hallucinations in agents’ reasoning, execution, perception, memory, and communication.
Continual Self-evolution Capacity.
@@
The analysis of agent hallucinations is typically conducted under the assumption of fixed user goals and static environments. However, in practice, both user demands and environmental configurations evolve continuously over time. To remain effective, agents must possess continual self-evolution capabilities that allow them to dynamically adapt to shifting goals and changing conditions. Throughout this process, agents can consistently refine their cognition to mitigate agent hallucinations stemming from outdated knowledge or delayed updates. Based on this important need, integrating the lifelong learning paradigm (Biesialska et al., 2020) with agents to endow them with more effective dynamic adaptation capabilities represents a promising solution.
Foundation Architecture Upgrade.
@@
Current LLM-based agents primarily rely on the Transformer architecture. However, this architecture faces challenges in handling long-context information and suffers from high computational complexity, which have gradually revealed performance bottlenecks and contributed to the emergence of hallucination issues. Recent works have explored more effective architectural upgrades, such as introducing linear-complexity modules (Dao et al., 2022) as Transformer components, integrating neural-symbolic systems (Cheng et al., 2025) to enhance model interpretability through symbolic reasoning, and leveraging automated machine learning (Baratchi et al., 2024) to design and compose optimal model architectures. In addition, agents typically require a predefined workflow to organize the execution of their various components. While this fixed pattern enhances systematization and controllability, it also reduces flexibility, resulting in limited adaptability and poor scalability. Designing a dynamic self-scheduling agentic system (Li et al., 2024d) that can autonomously organize task execution and coordinate multi-agent collaboration is a critical direction for future research.
6. Conclusion
@@
This paper presents a comprehensive survey of hallucination issues in LLM-based agents, with the goal of consolidating past progress, clarifying current challenges, and outlining future opportunities. We begin by distinguishing agent components into internal states and external behaviors, and, from this perspective, propose a taxonomy of hallucination types occurring at different stages. We then provide an in-depth overview of each type and identify eighteen underlying causes. Furthermore, we summarize ten general approaches for hallucination mitigation, together with corresponding detection methods. Finally, we discuss promising research directions to guide future exploration in this rapidly evolving domain. We believe this survey can serve as a valuable resource for researchers and practitioners, inspiring and facilitating further progress in the development of LLM-based agents.

### v2

@@ V Future Directions
1) Hallucinatory Accumulation Investigation. Most existing studies investigate hallucination instances and their underlying causes within a single agent loop. However, as we emphasize, agent decision-making is inherently a multi-step and sequential process, in which hallucinations can accumulate and amplify over time. In such cases, hallucinations may initially appear as minor issues, but their iterative accumulation can ultimately lead to severe consequences. Compared to single-step scenarios, hallucination accumulation presents a significantly more complex challenge. Addressing it requires a comprehensive analysis of the agent’s entire decision-making process, thereby facilitating early hallucination detection and mitigation.
@@
2) Accurate Hallucinatory Location. As previously mentioned, in contrast to hallucinations in traditional language models that are typically manifested in textual generation errors, agent hallucinations are far more complex, involving full-chain error propagation across multiple interdependent components. While hallucination taxonomy and attribution we presented in Section III have mapped these issues to specific modules within the agent, promptly and accurately locate the source of agent hallucinations in the final outputs remains a significant challenge. This difficulty stems from the fact that agent hallucinations may arise at any stage of the decision-making pipeline and often exhibit complex characteristics such as hallucinatory accumulation and inter-module dependency. In fact, the limited amount of research on agent hallucination detection in Section IV-B and the difficulty of conducting such works are also related to this issue. Therefore, future study should focus on designing agent systems capable of modeling and tracing their entire execution trajectories [zhang2025agent]. For example, lightweight checkpoints can be injected at each stage to verify whether hallucinations have occurred.
@@
3) Hallucination Mechanistic Interpretability. Mechanistic interpretability (MI) seeks to uncover how hidden representations and internal components of neural networks give rise to specific behaviors [olah2022mechinterp]. Therefore, MI provides a natural pathway for diagnosing and mitigating hallucinations at the root cause. Recent studies demonstrate that MI techniques, such as feature analysis [ji2025calibrating] and causal pathway tracing [jin2024cutting] can reveal the internal sources of hallucinations in LLMs and provide insights for mitigation. However, extending MI to LLM-based agents introduces new challenges. Unlike LLMs, where interpretability analysis is typically performed under single-step prediction with controlled prompt–response settings, agent hallucinations emerge through multi-step sequential interactions involving reasoning, tool use, memorization, and MAS communication. This dynamic process substantially complicates the controlled interventions required for mechanistic analysis. Future research should therefore adapt MI techniques to account for these dynamic and interconnected processes, enabling more precise diagnosis and systematic mitigation of agent hallucinations.
@@
4) Unified Benchmark Construction. Existing benchmarks for agent hallucinations are often limited to a specific hallucination type. For example, Guan et al. propose HALLUSIONBENCH to evaluate hallucination issues in visual-language reasoning [Guan_2024_CVPR]; Hu et al. propose MemoryAgentBench to evaluate hallucination issues in agents during memory retrieval and update [hu2025evaluating]; Zhang et al. introduce ToolBH to evaluate execution hallucinations in tool use across different scenarios from both depth and breadth perspectives [r55]. Therefore, there is a lack of a unified benchmark of hallucination evaluation that can define diverse hallucination scenarios and adopt various evaluation metrics to comprehensively assess the extent of hallucinations in agents’ reasoning, execution, perception, memory, and communication.
@@
5) Continual Self-evolution Capacity. The analysis of agent hallucinations is typically conducted under the assumption of fixed user goals and static environments. However, in practice, both user demands and environmental configurations evolve continuously over time. To remain effective, agents must possess continual self-evolution capabilities that allow them to dynamically adapt to shifting goals and changing conditions. Throughout this process, agents can consistently refine their cognition to mitigate agent hallucinations stemming from outdated knowledge or delayed updates. Based on this important need, integrating the lifelong learning paradigm [biesialska2020continual] with agents to endow them with more effective dynamic adaptation capabilities represents a promising solution.
@@
6) Foundation Architecture Upgrade. Current LLM-based agents primarily rely on the Transformer architecture. However, this architecture faces challenges in handling long-context information and suffers from high computational complexity, which have gradually revealed performance bottlenecks and contributed to the emergence of hallucination issues. Recent works have explored more effective architectural upgrades, such as introducing linear-complexity modules [dao2022flashattention] as Transformer components, integrating neural-symbolic systems [cheng2025neural] to enhance model interpretability through symbolic reasoning, and leveraging automated machine learning [baratchi2024automated] to design and compose optimal model architectures. In addition, agents typically require a predefined workflow to organize the execution of their various components. While this fixed pattern enhances systematization and controllability, it also reduces flexibility, resulting in limited adaptability and poor scalability. Designing a dynamic self-scheduling agentic system [li2024autoflow] that can autonomously organize task execution and coordinate multi-agent collaboration is a critical direction for future research.
VI Conclusion
@@
This paper presents a comprehensive survey of hallucination issues in LLM-based agents, with the goal of consolidating past progress, clarifying current challenges, and outlining future opportunities. We begin by distinguishing agent components into internal states and external behaviors, and, from this perspective, propose a taxonomy of hallucination types occurring at different stages. We then provide an in-depth overview of each type and identify seventeen underlying causes. Furthermore, we summarize ten general approaches for hallucination mitigation, together with corresponding detection methods. Finally, we discuss promising research directions to guide future exploration in this rapidly evolving domain. We believe this survey can serve as a valuable resource for researchers and practitioners, inspiring and facilitating further progress in the development of LLM-based agents.
Appendix A Loop of LLM-based Multi-agent System
Different from the single-agent setting, each LLM-based agent in the MAS must communicate with other agents to accomplish goals. To model these processes, a communication structure $G_{t}$ is introduced among $N$ agents. Furthermore, $G_{t}$ would evolve with time to align with the dynamic adjustments of MAS. Therefore, compared with Section II-B, the loop of LLM-based MAS includes two additional procedures: Broadcasting and Structure Evolution.
    $p_{t}^{i}=\Phi^{i}(b_{t}^{i},g).$ (15)
    $a_{t}^{i}=E^{i}(b_{t}^{i},p_{t}^{i}).$ (16)
    $\big\{q_{t}^{i\rightarrow j}\ \big|\ j\in N(i)\big\}=B(b_{t}^{i},p_{t}^{i},G_{t}).$ (17)
    $r_{t}=R\left(s_{t},\{a_{t}^{i}\}_{i=1}^{N}\right).$ (18)
    ${\rm Pr}\left(s_{t+1}\big|\ s_{t},\{a_{t}^{i}\}_{i=1}^{N}\right)=T\left(s_{t},\{a_{t}^{i}\}_{i=1}^{N}\right).$ (19)
    $o_{t+1}^{i}=Z^{i}\big(s_{t+1},b_{t}^{i},a_{t}^{i},\big\{q_{t}^{j\rightarrow i}\ \big|\ j\in N(i)\big\}\big),$ (20)
    $m_{t+1}^{i}=L_{M}^{i}\left(m_{t}^{i},a_{t}^{i},\big\{q_{t}^{j\rightarrow i}\ \big|\ j\in N(i)\big\},o_{t+1}^{i}\right).$ (21)
  • Belief Update: Then the agent $i$ refines its belief state as follows,
    $b_{t+1}^{i}=L_{B}^{i}\left(b_{t}^{i},m_{t+1}^{i},a_{t}^{i},\big\{q_{t}^{j\rightarrow i}\ \big|\ j\in N(i)\big\},r_{t}^{i},o_{t+1}^{i},g\right).$ (22)
    $G_{t+1}=U\left(G_{t},\{b_{t+1}^{i}\}_{i=1}^{N}\right).$ (23)
Figure 5: The depiction of different types of agent hallucinations, each illustrated with a representative example. The detailed explanation is given in Appendix B.
Appendix B Hallucination Example Explanation
As shown in Fig. 5, we present the representative examples of each type of agent hallucinations as follows,
  • Goal Understanding Hallucinations. In this example, the user asks the agent to “recommend a restaurant suitable for dining with elders”. The underlying goal is clearly to find a restaurant whose cuisine is light and easy to digest and whose environment is appropriate for family gatherings. However, the agent recommends “Spicy World”, a restaurant specializing in very spicy crayfish and boiled fish, thereby completely ignoring the prerequisite “suitable for elders”.
  • Intention Decomposition Hallucinations. In this example, the user asks the agent to plan a “family picnic for 10 people” and explicitly requires that the total budget for “purchasing food” must not exceed 500 yuan. Clearly, this plan must include a sub-intentions for “cost calculation” to verify whether the total expenditure satisfies the budget constraint. However, when decomposing the task, the agent only considers two steps: “choose a venue” and “buy food”, thereby omitting a crucial component in the decomposition of the user’s intention.
  • Planning Generation Hallucinations. This example illustrates hallucinations arising in the process of generating a concrete plan for a sub-intention derived from intention decomposition, namely “screen venues”. Given the user’s requirements, the chosen venue should explicitly allow picnicking. Nevertheless, the agent selects merely “a lawn that is open all day”, conflating “opening hours” with “time during which picnics are permitted”. In other words, the model hallucinates at the planning stage for the sub-intentions, failing to align the generated plan with the constraints specified in the user’s intent.
  • Perception Hallucinations. In this example, the user asks the agent to describe the content of an image. In the image, a chair is located to the right of a bookshelf, yet the agent responds that “the chair is to the left of the bookshelf”. This error is a prototypical perceptual hallucination, in which the agent’s visual understanding or spatial localization is systematically biased.
  • Memory Retrieval Hallucinations. Here, the user asks the agent to “tell me about tomorrow’s meeting schedule”. The memory store contains two relevant events: a project discussion meeting at 10:00 a.m. today and a customer meeting at 1:00 p.m. tomorrow. However, the agent fails to correctly anchor the events to their respective dates and instead treats both meetings as if they were scheduled for tomorrow. This illustrates that hallucinations can arise at the memory retrieval stage, where agents mis-retrieve or mis-align temporally indexed information.
  • Memory Update Hallucinations. In this example, the user instructs the agent to “cancel all meetings tomorrow morning”. During the memory update process, however, the agent erroneously clears the entire schedule for the following day, including the afternoon meeting. This reflects a hallucination in self-managed memory editing: the agent performs an overgeneralized or incorrect update, leading to a memory update hallucination.
  • Tool Selection Hallucinations. The user plans to travel to Kyoto in autumn and asks the agent to recommend tourist attractions. The agent incorrectly calls a non-existent tool named “get kyoto travel info”, whereas the correct tool in the system should be “recommend tourist spots”. This indicates that, when selecting external tools, the agent may “invent” plausible-sounding but non-existent APIs or function names, thereby exhibiting tool selection hallucinations.
  • Tool Calling Hallucinations. Similar to the previous case, this time the agent selects the correct tool, “recommend tourist spots”, but arbitrarily appends an extra parameter language=“English”, which is not part of the tool’s actual specification. This demonstrates that even when the tool name is correct, the agent may hallucinate at the parameter level—engaging in hallucinatory argument extension—which can cause the call to fail or yield unintended behavior.
  • Communication Hallucinations. This example illustrates an “echo chamber effect” in multi-agent communication, a particularly typical form of communication hallucination. Initially, the client states that they require a “code review”. As the message passes along a chain of agents, the first agent mishears it as “cold review”, the second relays it as “gold review”, and the third finally interprets it as “old review”. Through multiple rounds of transmission, the information is progressively distorted until it becomes completely detached from the original meaning. This hallucination shows that, when multiple agents collaborate, minor mis-hearings or mis-interpretations can be amplified, producing a “telephone game” style accumulation of deviations.